# Higgs Phases and Boundary Criticality in memoriam Peter Higgs

### v1

@@ IV.1.2 Open Boundaries and Global Matter Symmetry
where $\partial X_{k+1}$ denotes the set of $(k+1)$-cells in the boundary layer. As in Section II, this boundary action does not allow the matter field $\theta$ to tunnel through the boundary meaning that all the charged objects are closed and contained inside the bulk.
In the presence of these boundary conditions, the matter charge operators, Eq. 88, can generate a physical symmetry. For $\cramped{k>1}$ we must take care to consider how the surfaces $\tilde{\Sigma}$, which live in the dual lattice, can terminate at the boundary. Because we chose a “rough” boundary, as in Fig. 2, with a layer of cells sticking out from the edge of the system, the dual lattice boundary is flat. This is illustrated in Fig. 10(a) which shows the direct lattice bulk and boundary as in Fig. 2, along with the dual lattice in red and purple. The bulk part of the dual lattice is colored purple, while the boundary layer of the dual lattice is colored red and can be seen to form a flat surface without any protruding cells. The charge operators $\hat{Q}(\tilde{\Sigma})$ can terminate on this flat dual boundary layer.
For concreteness, consider the case $\cramped{k=2}$ in $\cramped{d=3}$, in which case the matter charges are strings and $\tilde{\Sigma}$ is a two-dimensional membrane in the dual lattice. Due to our boundary conditions the charge strings cannot terminate at the boundary, but the electric membranes can exit the system through the boundary. Referring to Fig. 10(b), consider a state with a single charged string wrapping around a periodic direction, as shown in Fig. 10(b) by the red line, attached to an electric membrane which exits through the boundary, illustrated by the yellow surface. This charged string is detected by taking the membrane $\tilde{\Sigma}$ to intersect it transversely, as illustrated by the cyan surface. The charge associated to the surface, Eq. 88, is then related to the electric flux through the boundary via the Gauss law,
$\hat{Q}(\tilde{\Sigma})=\langle-\mathrm{d}^{\dagger}\hat{E},\delta_{\tilde{\Sigma}}\rangle=-\langle\hat{E},\mathrm{d}\delta_{\tilde{\Sigma}}\rangle=-\langle\hat{E},\delta_{\partial\tilde{\Sigma}}\rangle\equiv\hat{Q}_{\text{bdry}}(\partial\tilde{\Sigma}).$ (94)
In the figure, $\partial\tilde{\Sigma}$ is shown as a dark green line which pierces a collection of boundary plaquettes (green squares with black borders). According to Eq. 94, the amount of charge measured by $\hat{Q}({\tilde{\Sigma}})$ is equal to the number of electric branes exiting through the boundary measured by the set of plaquettes pierced by $\partial\tilde{\Sigma}$. In summary, the physical matter symmetry is generated by charge operators supported on $\tilde{\Sigma}$ which terminate at the boundary and, by the Gauss law, acts on the gauge field $A$ on the boundary elements pierced by $\partial\tilde{\Sigma}$.
@@
Figure 11: An illustration of how the $\cramped{\kappa\to\infty}$ constraint $A=\mathrm{d}\theta$ turns closed membrane operators exiting through the boundary into open membrane operators terminating on a Higgs operator at the boundary. Each boundary $(k+1)$-cell $c$ has one bulk $k$-cell $c^{\prime}$ in its boundary. The constraint turns $A(c^{\prime})$ on this $k$-cell into $\mathrm{d}\theta(c^{\prime})=\theta(\partial c^{\prime})$. The end result is that all operators of the form $A(M)$ where $M$ exits the system through the boundary and is closed in the bulk are turned into open operators terminating on Higgs operators at the boundary, Eq. 96. We show (a,b) the case $k=1$ with a 0-form Higgs field, where (a) a string operator exiting the system at both ends (yellow), (b) turns into a pair of half-open string operators terminating on Higgs operators (red); and (c,d) the case $k=2$ with a 1-form Higgs field, where (c) a closed membrane operator exiting the system (yellow), (d) turns into a “half-open” membrane terminating on a Higgs string (red).
IV.1.3 Boundary Higher-Form Symmetry Breaking
Returning now to the higher-form gauge-Higgs action, Eq. 79, let us now see how the physical charge symmetry is spontaneously broken at the boundary. In the limit $\kappa\to\infty$ we have the constraint $A=\mathrm{d}\theta$ on every bulk $k$-cell. This completely freezes the bulk degrees of freedom, as is seen by rotating to unitary gauge, $\theta=0$, resulting in $\mathrm{d}A=0$. A covariant way to see this is that the gauge field operators $A(M)$ for $k$-dimensional closed surfaces $M$ trivialize—if $M$ is contained entirely within the bulk, then $\cramped{A(M)=\mathrm{d}\theta(M)=\theta(\partial M)=0}$ since $M$ is closed. This constraint is not imposed on the field variables on the cells touching the vacuum, however. As a result, if $M$ exist through the boundary, then we can decompose it into two pieces, $\cramped{M=M|_{X}+M|_{\partial X}}$, where $M|_{X}$ is the part of $M$ supported on bulk cells and $M|_{\partial X}$ is the part supported on boundary cells. Since $\cramped{\partial M=0}$, we have $\cramped{\partial M|_{\partial X}=-\partial M|_{X}}$. As a result, the gauge field operators decompose as
@@ IV.1.3 Boundary Higher-Form Symmetry Breaking
In other words, for closed surfaces $M$ exiting through the boundary, the $\cramped{\kappa\to\infty}$ constraint $A=\mathrm{d}\theta$ reduces $A(M)$ to a half-open Wilson operator terminating on Higgs operators as soon as it touches the bulk. Similarly, longer half-open Wilson surfaces ending on the Higgs field in the bulk, e.g. Eq. 16, are reduced to short half-open Wilson surfaces ending on the Higgs field at the boundary.
This is depicted in Fig. 11 for the cases $k=1$ and $k=2$, where $M$ is taken to be the smallest Wilson surfaces which are closed in the bulk, $M=\partial c$ for $c\in\partial X_{k+1}$. In the case $k=1$, depicted in Fig. 11(a), $c$ is an oriented boundary plaquette (green), and $\partial c$ consists of three links (yellow), one of which is in the bulk. We decompose $\partial c$ into the two links in the boundary, denoted $\partial c|_{\partial X}$, and the one link in the bulk, denoted $\partial c|_{X}$. The gauge field on the link in the bulk trivializes by the constraint $A=\mathrm{d}\theta$ into $\theta$ evaluated at its two endpoints, i.e. $\cramped{\theta(\partial(\partial c|_{X}))}$, depicted as red spheres in Fig. 11(b). These are the degrees of freedom appearing in the boundary action in Eq. 22. In the case $k=2$, Fig. 11(c), $c$ is a three-dimensional cube and $\partial c$ is a set of five plaquettes (yellow). In Fig. 11(d), it decomposes into four plaquettes in the boundary (yellow), $\partial c|_{\partial X}$, and one plaquette in the bulk, $\partial c|_{X}$. The constraint turns the gauge field on this bulk plaquette into a Higgs string operator on its boundary (red). The extension to larger $k$ is obvious, but can’t be illustrated.
Using this, we can see that the boundary action Eq. 93 contains precisely these minimal operators. In the general case, the action reduces in this limit to
@@
We have already seen in Section II how in the case $k=1$ this can be recast as a 0-form XY model, Eq. 22, which we can now extend to the case $k>1$. Note that every boundary $k$-cell is associated to a unique bulk $(k-1)$-cell (the one which trivializes under the $A=\mathrm{d}\theta$ constraint). We define composite degrees of freedom for each such pair,
@@
where $A_{\partial X}(c)$ is $A$ evaluated on the unique boundary cell corresponding to $c$, which we treat as a $\cramped{(k-1)}$-form rather than a $k$-form. Note that $\vartheta$ are not gauge invariant when $\cramped{k>1}$, because they would create open charged objects, but we can combine them to construct the gauge-invariant degrees of freedom appearing in Eq. 97. For example, in the case $\cramped{k=2}$, each $\vartheta$ consists of $A$ on a boundary plaquette and $-\theta$ on the bulk link it touches, four of which combine to form the gauge-invariant composite object in Fig. 11(d). Let us denote the layer of bulk cells which touch the boundary layer, consisting of cells up to dimension $\cramped{d-1}$, as $Y$, illustrated in Fig. 10 by the dark layer between $X$ and $\partial X$. Each $(k+1)$-cell $c\in\partial X$ is associated to a unique $k$-cell $c_{Y}\in Y$. We treat $\vartheta$ as a $\cramped{(k-1)}$-form gauge field in $Y$, so that its exterior derivative is given by
@@
where we identified $\partial c_{Y}$ with $\partial c|_{\partial X}$. We can then write the boundary action Eq. 97 as
@@
Equation 100 is precisely the action of a $(k-1)$-form U(1) gauge theory defined at the boundary of the system in terms of composite half-open Wilson operators. In the case $k=1$ it reduces to the 0-form XY model identified in Section II.
We conclude that a $k$-form Abelian-Higgs model in $D$ dimensions Higgses in the $\kappa\to\infty$ limit down to a $(k-1)$-form gauge theory on the $\cramped{D-1}$ dimensional boundary. If $\cramped{D>k+2}$ ($\cramped{k_{\text{m}}>0}$), this implies there will be a boundary phase transition. For $k=1$ this will be in the $(D-1)$-dimensional XY universality class, with a symmetry-broken phase at large $\beta$. For $k>1$ it will be a confinement-deconfinement transition (expected to be first-order [61]), with a confined phase at small $\beta$ and a deconfined phase at large $\beta$. This reduces to the results reported in Section II for $k=1$ and $D=4$. The symmetry that is spontaneously broken at large $\beta$ is the $(k-1)$-form matter symmetry, corresponding to the electric symmetry of the boundary gauge theory when $k>1$, under which $\mathrm{d}\vartheta$ (as a half-open Wilson operator) is charged. A summary of this result is given in the table Table 1 for $k\leq 3$ and $2\leq D\leq 6$, which may be extended straightforwardly to all $k$ and $D$. In the cases where $\cramped{D=k+2}$ ($\cramped{k_{\text{m}}=0}$), a generalized Mermin-Wagner theorem [4, 6] (equivalently, the Polyakov mechanism or magnetic instanton proliferation [58]) will prevent the symmetry from breaking.
Table 1: The boundary theories of $k$-form U(1) Abelian-Higgs models in $D$ spacetime dimensions in the $\cramped{\kappa\to\infty}$ limit and their phase transitions. The pattern is that a $k$-form Abelian Higgs model Higgses down to a $(k-1)$-form gauge theory on the boundary without matter. The number $\cramped{k_{\mathrm{m}}=D-(k+2)}$ indicates the dimension of the magnetic worldlines of the bulk theory, $\tilde{j}_{\mathrm{m}}$ in Eq. 107. Note that in the $\beta\to\infty$ limit the bulk has a corresponding $k_{\mathrm{m}}$-form symmetry. Gray boxes indicates cases where the gauge field has no dynamics. The top non-trivial box of each column, which has $\cramped{D=k+2}$ ($\cramped{k_{\mathrm{m}}=0}$), are affected by magnetic instanton proliferation (the Polyakov mechanism or a generalized Mermin-Wagner theorem forbidding boundary symmetry breaking) and do not exhibit a boundary phase transition, except in the case $\cramped{k=1}$ in $\cramped{D=3}$, which can exhibit a BKT transition.
IV.2 Electric-Magnetic Dual Picture
@@
IV.2.1 Open Boundary Duality Transformation
@@
We interpret the $k$-form $j_{\mathrm{e}}$ as the worldlines swept out by the $(k-1)$-dimensional Higgs excitations (carrying integer electric charge), and $e$ as the worldsheets swept out by the $k$-form electric field. The first delta function is just the Gauss law—it tells us that the worldsheets of electric flux must terminate on the worldlines of the Higgs charges. The second delta function enforces that the the worldlines $j_{\mathrm{e}}$ are “divergence-free,” i.e. they form closed $k$-dimensional surfaces, corresponding to global conservation of charge. Note that, because the electric worldsheets can end, the electric 1-form symmetry $\mathrm{d}^{\dagger}e=0$, present when $\kappa=0$, is explicitly broken, though it can be restored as an emergent symmetry at energies below the charge gap.
With the choice of “electric” boundary conditions (Fig. 2), we have the constraint $j_{\mathrm{e}}=0$ on all links extending from the bulk to the vacuum in Eq. 103, because there is no $\mathrm{d}\theta$ term for these links, while $e$ can be non-zero on the boundary plaquettes. This means that electric charge cannot enter or exit the system, but electric flux can. The fluxes in the bulk form closed surfaces unless they terminate on charges. In the presence of the boundary these fluxes may be cut off at the boundary without terminating on charges.
@@
to approximate the partition function by999 We use the shorthand $(\omega)^{2}$ to denote $\langle\omega,\omega\rangle$.
@@
coupled respectively to integer currents $b$ and $j_{\text{m}}$ via Poisson resummation. Lastly, we move to the dual lattice, replacing $\mathrm{d}^{\dagger}\alpha\leftrightarrow\mathrm{d}\tilde{\alpha}$ for each $p$-form $\alpha$, where $\tilde{\alpha}$ is a ($D-p$)-form in the dual lattice, the discrete Hodge dual of $\alpha$. We thus obtain the dual partition function
@@
where $\tilde{j}_{\mathrm{m}}$ and $\tilde{A}_{\mathrm{m}}$ are $(\cramped{D-(k+2)})$ forms, while $\tilde{b}$ and $\tilde{H}$ are $(\cramped{D-(k+1)})$ forms, with the constraint that $\mathrm{d}\tilde{H}$ is zero within the dual boundary layer coming from the delta function in Eq. 105. Finally, let us rescale the fields by absorbing the $2\pi$ into the definition of $\tilde{A}_{\mathrm{m}}$ and $\tilde{H}$, to obtain the dual action
@@
$\delta S_{\text{dual}}^{\text{bulk}}=-i[\langle\lambda,\tilde{j}_{\mathrm{m}}\rangle-\langle\mathrm{d}\tilde{\lambda},\tilde{b}\rangle]=-i\langle\tilde{\lambda},\tilde{j}_{\mathrm{m}}-\mathrm{d}^{\dagger}\tilde{b}\rangle.$ (109)
@@
This says that the magnetic monopole worldlines form the boundaries of the Higgs vortex worldsheets. This is precisely the magnetic Gauss law, i.e. it says that magnetic monopoles are the sources of magnetic strings (compare to the electric Gauss law Eq. 103). This is familiar from superconductor phenomenology—vortex cores carry magnetic flux. It also follows from this that $\mathrm{d}^{\dagger}\tilde{j}_{\mathrm{m}}=0$, i.e. that the magnetic worldlines are closed and magnetic charge is conserved, which follows from the “gauge-of-gauge” invariance $\cramped{\tilde{A}_{\mathrm{m}}\to\tilde{A}_{\mathrm{m}}+\mathrm{d}\tilde{\alpha}}$ for arbitrary $\tilde{\alpha}$.
This theory can be recast as a U(1) gauge theory as follows. Summing over $\tilde{b}$ undoes one of the Poisson resummations and forces $\tilde{H}=2\pi\tilde{h}$, where the integer field $\tilde{h}$ is the Hodge dual of $h$ in Eq. 105. The residual gauge symmetry is $\cramped{\tilde{A}_{\mathrm{m}}\to\tilde{A}_{\mathrm{m}}+2\pi\tilde{l}}$ and $\cramped{\tilde{H}\to\tilde{H}-2\pi\mathrm{d}\tilde{l}}$, where $\tilde{l}$ is an integer shift, meaning that the gauge-invariant configuration space for $\tilde{A}_{\mathrm{m}}$ is actually $\mathbbm{R}/2\pi\mathbbm{Z}$ and for $\tilde{H}$ is $2\pi\mathbbm{Z}$. The resulting theory is therefore a Villainized $\cramped{k_{\text{m}}}$-form U(1) gauge theory [62, 20] (Eq. 84) coupled to magnetic currents, the dual of the original Abelian Higgs model model, Eq. 101, which was coupled to electric currents. The integer gauge field $\tilde{h}$ measures the winding numbers of the compact $\tilde{A}_{\mathrm{m}}$, and its fluxes $\mathrm{d}\tilde{H}$ are the homotopy defects of $\tilde{A}_{\mathrm{m}}$, which are the electric charges of the original theory. They act as sources for the fluxes $\mathrm{d}\tilde{A}_{\mathrm{m}}$, which are the electric strings in the direct lattice.

### v2

@@ 4.1.2 Open Boundaries and Global Matter Symmetry
where $\partial X_{k+1}$ denotes the set of $(k+1)$-cells in the boundary layer. As in Section˜2, this boundary action does not allow the matter field $\theta$ to tunnel through the boundary meaning that all the charged objects are closed and contained inside the bulk.
[…]
For concreteness, consider the case $\cramped{k=2}$ in $\cramped{d=3}$, in which case the matter charges are strings and $\tilde{\Sigma}$ is a two-dimensional membrane in the dual lattice. Due to our boundary conditions the charge strings cannot terminate at the boundary, but the electric membranes can exit the system through the boundary. Referring to Fig.˜11(b), consider a state with a single charged string wrapping around a periodic direction, as shown in Fig. 11(b) by the red line, attached to an electric membrane which exits through the boundary, illustrated by the yellow surface. This charged string is detected by taking the membrane $\tilde{\Sigma}$ to intersect it transversely, as illustrated by the cyan surface. The charge associated to the surface, Eq.˜90, is then related to the electric flux through the boundary via the Gauss law,
[…]
@@
Figure 12: An illustration of how the $\cramped{\kappa\to\infty}$ constraint $A=\mathrm{d}\theta$ turns closed membrane operators exiting through the boundary into open membrane operators terminating on a Higgs operator at the boundary. Each boundary $(k+1)$-cell $c$ has one bulk $k$-cell $c^{\prime}$ in its boundary. The constraint turns $A(c^{\prime})$ on this $k$-cell into $\mathrm{d}\theta(c^{\prime})=\theta(\partial c^{\prime})$. The end result is that all operators of the form $A(M)$ where $M$ exits the system through the boundary and is closed in the bulk are turned into open operators terminating on Higgs operators at the boundary, Eq.˜98. We show (a,b) the case $k=1$ with a 0-form Higgs field, where (a) a string operator exiting the system at both ends (yellow), (b) turns into a pair of half-open string operators terminating on Higgs operators (red); and (c,d) the case $k=2$ with a 1-form Higgs field, where (c) a closed membrane operator exiting the system (yellow), (d) turns into a “half-open” membrane terminating on a Higgs string (red).
4.1.3 Boundary Symmetry Breaking
Returning now to the higher-form gauge-Higgs action, Eq.˜81, let us now see how the physical charge symmetry is spontaneously broken at the boundary. In the limit $\kappa\to\infty$ we have the constraint $A=\mathrm{d}\theta$ on every bulk $k$-cell. This completely freezes the bulk degrees of freedom, as is seen by rotating to unitary gauge, $\theta=0$, resulting in $\mathrm{d}A=0$. A covariant way to see this is that the gauge field operators $A(M)$ for $k$-dimensional closed surfaces $M$ trivialize—if $M$ is contained entirely within the bulk, then $\cramped{A(M)=\mathrm{d}\theta(M)=\theta(\partial M)=0}$ since $M$ is closed. This constraint is not imposed on the field variables on the cells touching the vacuum, however. As a result, if $M$ exist through the boundary, then we can decompose it into two pieces, $\cramped{M=M|_{X}+M|_{\partial X}}$, where $M|_{X}$ is the part of $M$ supported on bulk cells and $M|_{\partial X}$ is the part supported on boundary cells. Since $\cramped{\partial M=0}$, we have $\cramped{\partial M|_{\partial X}=-\partial M|_{X}}$. As a result, the gauge field operators decompose as
@@ 4.1.3 Boundary Symmetry Breaking
In other words, for closed surfaces $M$ exiting through the boundary, the $\cramped{\kappa\to\infty}$ constraint $A=\mathrm{d}\theta$ reduces $A(M)$ to a half-open Wilson operator terminating on Higgs operators as soon as it touches the bulk. Similarly, longer half-open Wilson surfaces ending on the Higgs field in the bulk, e.g. Eq.˜16, are reduced to short half-open Wilson surfaces ending on the Higgs field at the boundary.
This is depicted in Fig.˜12 for the cases $k=1$ and $k=2$, where $M$ is taken to be the smallest Wilson surfaces which are closed in the bulk, $M=\partial c$ for $c\in\partial X_{k+1}$. In the case $k=1$, depicted in Fig.˜12(a), $c$ is an oriented boundary plaquette (green), and $\partial c$ consists of three links (yellow), one of which is in the bulk. We decompose $\partial c$ into the two links in the boundary, denoted $\partial c|_{\partial X}$, and the one link in the bulk, denoted $\partial c|_{X}$. The gauge field on the link in the bulk trivializes by the constraint $A=\mathrm{d}\theta$ into $\theta$ evaluated at its two endpoints, i.e. $\cramped{\theta(\partial(\partial c|_{X}))}$, depicted as red spheres in Fig.˜12(b). These are the degrees of freedom appearing in the boundary action in Eq.˜22. In the case $k=2$, Fig.˜12(c), $c$ is a three-dimensional cube and $\partial c$ is a set of five plaquettes (yellow). In Fig.˜12(d), it decomposes into four plaquettes in the boundary (yellow), $\partial c|_{\partial X}$, and one plaquette in the bulk, $\partial c|_{X}$. The constraint turns the gauge field on this bulk plaquette into a Higgs string operator on its boundary (red). The extension to larger $k$ is obvious, but can’t be illustrated.
Using this, we can see that the boundary action Eq.˜95 contains precisely these minimal operators. In the general case, the action reduces in this limit to
@@
We have already seen in Section˜2 how in the case $k=1$ this can be recast as a 0-form XY model, Eq.˜22, which we can now extend to the case $k>1$. Note that every boundary $k$-cell is associated to a unique bulk (k1)(k-1)-cell (the one which trivializes under the $A=\mathrm{d}\theta$ constraint). We define composite degrees of freedom for each such pair,
@@
where $A_{\partial X}(c)$ is $A$ evaluated on the unique boundary cell corresponding to $c$, which we treat as a $\cramped{(k-1)}$-form rather than a $k$-form. Note that $\vartheta$ are not gauge invariant when $\cramped{k>1}$, because they would create open charged objects, but we can combine them to construct the gauge-invariant degrees of freedom appearing in Eq.˜99. For example, in the case $\cramped{k=2}$, each $\vartheta$ consists of $A$ on a boundary plaquette and $\theta$ on the bulk link it touches, four of which combine to form the gauge-invariant composite object in Fig.˜12(d). Let us denote the layer of bulk cells which touch the boundary layer, consisting of cells up to dimension $\cramped{d-1}$, as $Y$, illustrated in Fig.˜11 by the dark layer between $X$ and $\partial X$. Each $(k+1)$-cell $c\in\partial X$ is associated to a unique $k$-cell $c_{Y}\in Y$. We treat $\vartheta$ as a $\cramped{(k-1)}$-form gauge field in $Y$, so that its exterior derivative is given by
@@
where we identified $\partial c_{Y}$ with $\partial c|_{\partial X}$. We can then write the boundary action Eq.˜99 as
@@
Equation˜102 is precisely the action of a (k1)(k-1)-form U(1) gauge theory defined at the boundary of the system in terms of composite half-open Wilson operators. In the case $k=1$ it reduces to the 0-form XY model identified in Section˜2.
We conclude that a $k$-form Abelian-Higgs model in $D$ dimensions Higgses in the $\kappa\to\infty$ limit down to a (k1)(k-1)-form gauge theory on the $\cramped{D-1}$ dimensional boundary. If $\cramped{D>k+2}$ ($\cramped{k_{\text{m}}>0}$), this implies there will be a boundary phase transition. For $k=1$ this will be in the (D1)(D-1)-dimensional XY universality class, with a symmetry-broken phase at large $\beta$. For $k>1$ it will be a confinement-deconfinement transition (expected to be first-order [7]), with a confined phase at small $\beta$ and a deconfined phase at large $\beta$. This reduces to the results reported in Section˜2 for $k=1$ and $D=4$. The symmetry that is spontaneously broken at large $\beta$ is the (k1)(k-1)-form matter symmetry, corresponding to the electric symmetry of the boundary gauge theory when $k>1$, under which $\mathrm{d}\vartheta$ (as a half-open Wilson operator) is charged. A summary of this result is given in the table Table˜1 for $k\leq 3$ and $2\leq D\leq 6$, which may be extended straightforwardly to all $k$ and $D$. In the cases where $\cramped{D=k+2}$ ($\cramped{k_{\text{m}}=0}$), a generalized Mermin-Wagner theorem [5, 8] (equivalently, the Polyakov mechanism or magnetic instanton proliferation [66]) will prevent the symmetry from breaking. This may imply that the Higgs regimes of these marginal cases, which do not exhibit bulk higher-form symmetry breaking (deconfinement), are qualitatively distinct from the cases exhibiting boundary symmetry breaking.
4.2 Electric-Magnetic Dual Picture
@@
4.2.1 Open Boundary Duality Transformation
@@
We interpret the $k$-form $j_{\mathrm{e}}$ as the worldlines swept out by the (k1)(k-1)-dimensional Higgs excitations (carrying integer electric charge), and $e$ as the worldsheets swept out by the $k$-form electric field. The first delta function is just the Gauss law—it tells us that the worldsheets of electric flux must terminate on the worldlines of the Higgs charges. The second delta function enforces that the the worldlines $j_{\mathrm{e}}$ are “divergence-free,” i.e. they form closed $k$-dimensional surfaces, corresponding to global conservation of charge. Note that, because the electric worldsheets can end, the electric 1-form symmetry $\mathrm{d}^{\dagger}e=0$, present when $\kappa=0$, is explicitly broken, though it can be restored as an emergent symmetry at energies below the charge gap.
With the choice of “electric” boundary conditions (Fig.˜2), we have the constraint $j_{\mathrm{e}}=0$ on all links extending from the bulk to the vacuum in Eq.˜105, because there is no $\mathrm{d}\theta$ term for these links, while $e$ can be non-zero on the boundary plaquettes. This means that electric charge cannot enter or exit the system, but electric flux can. The fluxes in the bulk form closed surfaces unless they terminate on charges. In the presence of the boundary these fluxes may be cut off at the boundary without terminating on charges.
@@
to approximate the partition function by101010 We use the shorthand $(\omega)^{2}$ to denote $\langle\omega,\omega\rangle$.
@@
coupled respectively to integer currents $b$ and $j_{\text{m}}$ via Poisson resummation. Lastly, we move to the dual lattice, replacing $\mathrm{d}^{\dagger}\alpha\leftrightarrow\mathrm{d}\tilde{\alpha}$ for each $p$-form $\alpha$, where $\tilde{\alpha}$ is a (DpD-p)-form in the dual lattice, the discrete Hodge dual of $\alpha$ [34]. We thus obtain the dual partition function
@@
where $\tilde{j}_{\mathrm{m}}$ and $\tilde{A}_{\mathrm{m}}$ are $(\cramped{D-(k+2)})$-forms, while $\tilde{b}$ and $\tilde{H}$ are $(\cramped{D-(k+1)})$-forms, with the constraint that $\mathrm{d}\tilde{H}$ is zero within the dual boundary layer coming from the delta function in Eq.˜107. Finally, let us rescale the fields by absorbing the $2\pi$ into the definition of $\tilde{A}_{\mathrm{m}}$ and $\tilde{H}$, to obtain the dual action
@@
$\delta S_{\text{dual}}^{\text{bulk}}=-i[\langle\tilde{\lambda},\tilde{j}_{\mathrm{m}}\rangle-\langle\mathrm{d}\tilde{\lambda},\tilde{b}\rangle]=-i\langle\tilde{\lambda},\tilde{j}_{\mathrm{m}}-\mathrm{d}^{\dagger}\tilde{b}\rangle$. (111)
@@
This says that the magnetic monopole worldlines form the boundaries of the Higgs vortex worldsheets. This is precisely the magnetic Gauss law, i.e. it says that magnetic monopoles are the sources of magnetic strings (compare to the electric Gauss law Eq.˜105). This is familiar from superconductor phenomenology—vortex cores carry magnetic flux. It also follows from this that $\mathrm{d}^{\dagger}\tilde{j}_{\mathrm{m}}=0$, i.e. that the magnetic charge worldlines are closed and magnetic charge is conserved, which follows from the “gauge-of-gauge” invariance $\cramped{\tilde{A}_{\mathrm{m}}\to\tilde{A}_{\mathrm{m}}+\mathrm{d}\tilde{\alpha}}$ for arbitrary $\tilde{\alpha}$.
Table 1: The boundary theories of $k$-form U(1) Abelian-Higgs models in $D$ spacetime dimensions in the $\cramped{\kappa\to\infty}$ limit and their phase transitions. The pattern is that a $k$-form Abelian Higgs model Higgses down to a (k1)(k-1)-form gauge theory on the boundary without matter. The number $\cramped{k_{\mathrm{m}}=D-(k+2)}$ indicates the dimension of the magnetic worldlines of the bulk theory, $\tilde{j}_{\mathrm{m}}$ in Eq.˜109. Note that in the $\cramped{\beta\to\infty}$ limit the bulk has a corresponding $k_{\mathrm{m}}$-form symmetry. Gray boxes indicates cases where the gauge field has no dynamics. The top non-trivial box of each column, which has $\cramped{D=k+2}$ ($\cramped{k_{\mathrm{m}}=0}$), are affected by magnetic instanton proliferation (the Polyakov mechanism or a generalized Mermin-Wagner theorem forbidding boundary symmetry breaking) and do not exhibit a boundary phase transition, except in the case $\cramped{k=1}$ in $\cramped{D=3}$, which can exhibit a BKT transition.
This theory can be recast as a U(1) gauge theory as follows. Summing over $\tilde{b}$ undoes one of the Poisson resummations and forces $\tilde{H}=2\pi\tilde{h}$, where the integer field $\tilde{h}$ is the Hodge dual of $h$ in Eq.˜107. The residual gauge symmetry is $\cramped{\tilde{A}_{\mathrm{m}}\to\tilde{A}_{\mathrm{m}}+2\pi\tilde{l}}$ and $\cramped{\tilde{H}\to\tilde{H}-2\pi\mathrm{d}\tilde{l}}$, where $\tilde{l}$ is an integer shift, meaning that the gauge-invariant configuration space for $\tilde{A}_{\mathrm{m}}$ is actually $\mathbbm{R}/2\pi\mathbbm{Z}$ and for $\tilde{H}$ is $2\pi\mathbbm{Z}$. The resulting theory is therefore a Villainized $\cramped{k_{\text{m}}}$-form U(1) gauge theory [69, 23] (Eq.˜86) coupled to magnetic currents, the dual of the original Abelian Higgs model model, Eq.˜103, which was coupled to electric currents. The integer gauge field $\tilde{h}$ measures the winding numbers of the compact $\tilde{A}_{\mathrm{m}}$, and its fluxes $\mathrm{d}\tilde{H}$ are the homotopy defects of $\tilde{A}_{\mathrm{m}}$, which are the electric charges of the original theory. They act as sources for the fluxes $\mathrm{d}\tilde{A}_{\mathrm{m}}$, which are the electric strings in the direct lattice.